# Correlation-Aware Graph Convolutional Networks for Multi-Label Node Classification

### v2

@@ 1. Introduction
Inspired by the success of multi-label learning on non-relational data (Liu et al., 2017; Wu et al., 2020; Wang et al., 2020), such as images and text, there has been a growing interest in applying multi-label learning to relational data, including graphs (Gao et al., 2019; Xiao et al., 2022b; Zhou et al., 2021). Among the few existing methods for multi-label node classification, most employ GCNs to map nodes to low-dimensional representations. They then follow the traditional multi-label learning paradigm to model the relationships between nodes and labels, as well as the inter-label relationships in the representation space. However, this paradigm overlooks the characteristics of graph data in multi-label scenarios, as Figure 1-(a), leading to challenges for GCNs. (1) Ambiguous Feature: In the single-label setting, nodes can aggregate patterns specific to a particular type of label from one neighborhood node by transforming the features. However, in the multi-label setting, the feature of one node may be related to multiple labels and the patterns represented by the feature are ambiguous. As a result, aggregating information from such ambiguous features will significantly impair the discriminative ability of the representation. Figure 1-(b) illustrates that nodes that allocate similar features may share a diverse number of labels in real-world graphs. (2) Ambiguous Topology: In the single-label setting, connected nodes typically share the same single label (also known as the homophily assumption), so the pattern propagated along the edge is usually deterministic. However, in the multi-label setting, both the connected nodes have multiple labels, and the patterns propagated along the edge are often ambiguous. This ambiguity makes it difficult to determine from which connected nodes we should aggregate specific label information. Aggregating information directly from all neighbors would further accumulate ambiguity, compromising the discriminative power of the representations learned by GCNs and eventually impacting the inferring of specific labels. Figure 1-(c) illustrates that connected nodes may share diverse numbers of labels in real-world graphs. It is essential to reduce the ambiguity so that the potential of GCNs in multi-label node classification can be fully released.
@@ 5.1.3. Implementation Setting
For all models, the embedding size is fixed to 64 for fair node classification. The batch size is set to 1024 for all models and the Adam optimizer is used. Experiments are conducted on the Centos system with NVIDIA RTX-3090 GPUs. The detailed experiment implementations of CorGCN are posed in Appendix B.3. Source code is available at https://github.com/YuanchenBei/CorGCN.
@@ 5.3. Ablation Study (RQ2)
  • Effectiveness of the correlation decomposed graph. The severe performance decline of w/o CGD, along with the diminished efficacy of w/o CFD and w/o CSD, demonstrates that the original graph topology and node features are suboptimal to multi-label node classification. The proposed correlation decomposed graph (CDG) with both correlation decomposed node features and topology structure can effectively benefit the classification.
@@ 5.6. Efficiency Study (RQ5)
To further evaluate whether CorGCN can balance the effectiveness and efficiency, we conduct the efficiency study of CorGCN compared with the top-3 performed baselines on both training and inference stage in Table 4.
From the results, we can find that our CorGCN can generally balance effectiveness and efficiency. For the training stage, the end-to-end label correlation modeling strategy employed by CorGCN yields an efficiency that is second to that of pure GCN. The negative sampling process costs additional time for ML-GCN during training. Then, during the inference phase, CorGCN achieves computational efficiency that is lower than top-3 performing baselines due to the multi-label structure decomposition while demonstrating a good enhancement in performance compared to these models.
Table 4. Efficiency study results on Humloc and PCG datasets each epoch in seconds.
@@ Appendix A Methodology Details
A.1. Analysis of the Objective Function on your Multi-Label Estimator
@@
Analysis of the effectiveness of mutual information estimator $\phi$ for multi-label correlation modeling can be found in previous works (Bai et al., 2022; Zhao et al., 2021). Then, The likelihood maximizing decoder $\theta$ can be utilized as a likelihood estimator of node feature and label prototype with the entropy KL divergence with focal enhancement (Lin et al., 2017) for target labels as follows:
@@ A.2. Complexity Analysis
The overall complexity of CorGCN arises from the following stages. During the correlation-aware graph decomposition, the calculation involves the projection process, incurring a computational cost of $\Theta(K\cdot n^{2})$. For the Correlation-Enhanced graph convolution, we postulate that the message passing process within a graph comprising $n$ nodes entails a cost of $\Theta_{\text{mp}}(n)$, which varies according to the specific GCN backbone employed. During the intra-label message passing, due to the process can be paralleled across $K$ label views, the overall expense of message passing is quantified as $\Theta_{\text{mp}}(n)$. Then, the cost associated with inter-label correlation propagation is $\Theta(K^{2}\cdot n)$. Thus, the overall time complexity of the process can be expressed as $\mathbf{T}(n)=\Theta(K(n^{2}+\Theta_{\text{mp}}(n)+K\cdot n))$.
As outlined in Sec. 4.4, this complexity can be reduced to $\mathbf{T}(n)=\Theta(K^{\prime}(n^{2}+\Theta_{\text{mp}}(n)+K^{\prime}n))$ for the large label space, where $K^{\prime}\ll K$, denoting a significant reduction in computational cost. In practice, the entire process can be batch-processed, where the similarity is calculated among the nodes within each batch, thus making the time complexity of each batch $B$ to $\mathbf{T}(B)=\Theta(K^{\prime}(B^{2}+\Theta_{\text{mp}}(B)+K^{\prime}B))$.
@@ A.3. Pseudocode of CorGCN
The overall pipeline of CorGCN includes (1) correlation-aware graph learning with feature and topology decompositions, and (2) correlation-aware graph convolution with intra-label message passing and inter-label correlation propagation. The general process of this pipeline is described as the pseudocode in Algorithm 1.

### v3

@@ 1. Introduction
Inspired by the success of multi-label learning on non-relational data (Liu et al., 2017; Wu et al., 2020; Wang et al., 2020), such as images and text, there has been a growing interest in applying multi-label learning to relational data, including graphs (Gao et al., 2019; Xiao et al., 2022b; Zhou et al., 2021). Among the few existing methods for multi-label node classification, most employ GCNs to map nodes to low-dimensional representations. They then follow the traditional multi-label learning paradigm to model the relationships between nodes and labels, as well as the inter-label relationships in the representation space. However, this paradigm overlooks the characteristics of graph data in multi-label scenarios, as Figure 1-(a), leading to challenges for GCNs. (1) Ambiguous Feature: In the single-label setting, nodes can aggregate patterns specific to a particular type of label from one neighborhood node by transforming the features. However, in the multi-label setting, the feature of one node may be related to multiple labels and the patterns represented by the feature are ambiguous. As a result, aggregating information from such ambiguous features will significantly impair the discriminative ability of the representation. Figure 1-(b) illustrates that nodes that allocate similar features may share a diverse number of labels in real-world graphs. (2) Ambiguous Topology: In the single-label setting, connected nodes typically share the same single label (also known as the edge homophily assumption), so the pattern propagated along the edge is usually deterministic. However, in the multi-label setting, both the connected nodes have multiple labels, and the patterns propagated along the edge are often ambiguous. This ambiguity makes it difficult to determine from which connected nodes we should aggregate specific label information. Aggregating information directly from all neighbors would further accumulate ambiguity, compromising the discriminative power of the representations learned by GCNs and eventually impacting the inferring of specific labels. Figure 1-(c) illustrates that edge-connected nodes may share diverse numbers of labels in real-world graphs. It is essential to reduce the ambiguity so that the potential of GCNs in multi-label node classification can be fully released.
@@ 5.1.3. Implementation Setting
For all models, the embedding size is fixed to 64 for fair node classification. The batch size is set to 1024 for all models and the Adam optimizer is used. Experiments are conducted on the Centos system with NVIDIA RTX-3090 GPUs. The detailed experiment implementations of CorGCN are posed in Appendix B.3 and the source code of CorGCN is available at the Github link.
@@ 5.3. Ablation Study (RQ2)
  • Effectiveness of the correlation decomposed graph. The severe performance decline of w/o CGD, along with the diminished efficacy of w/o CFD and w/o CSD, demonstrates that the original graph topology and node features are suboptimal for multi-label node classification. The proposed correlation decomposed graph (CDG) with both correlation decomposed node features and topology structure can effectively benefit the classification.
@@ 5.6. Efficiency Study (RQ5)
To further evaluate whether CorGCN can balance the effectiveness and efficiency, we conduct the efficiency study of CorGCN compared with the top-3 performed baselines on both the training and inference stages in Table 4.
From the results, we can find that our CorGCN can generally balance effectiveness and efficiency. For the training stage, the end-to-end label correlation modeling strategy employed by CorGCN yields an efficiency that is second to that of the original GCN. The negative sampling process costs additional time for ML-GCN during training. Then, during the inference phase, CorGCN has computational efficiency that is lower than top-3 performing baselines due to the multi-label structure decomposition, while demonstrating a good enhancement in performance compared to these models.
Table 4. Efficiency study results on Humloc and PCG datasets each convolution epoch in seconds.
@@ Appendix A Methodology Details
A.1. Analysis of the Objective Function on the Multi-Label Estimator
@@
Analysis of the effectiveness of the mutual information estimator $\phi$ for multi-label correlation modeling can be found in previous works (Bai et al., 2022; Zhao et al., 2021). Then, the likelihood maximizing decoder $\theta$ can be utilized as a likelihood estimator of node feature and label prototype with the entropy KL divergence with focal enhancement (Lin et al., 2017) for target labels as follows:
@@ A.2. Complexity Analysis
During the correlation-aware graph decomposition, the calculation involves the projection process, incurring a computational cost of $\Theta(K\cdot n^{2})$. For the Correlation-Enhanced graph convolution, we postulate that the message passing process within a graph comprising $n$ nodes entails a cost of $\Theta_{\text{mp}}(n)$, which varies according to the specific GCN backbone employed. During the intra-label message passing, since the process can be parallelized across $K$ label views, the overall expense of message passing is quantified as $\Theta_{\text{mp}}(n)$. Then, the cost associated with inter-label correlation propagation is $\Theta(K^{2}\cdot n)$. Thus, the overall time complexity of the process can be expressed as $\mathbf{T}(n)=\Theta(K(n^{2}+\Theta_{\text{mp}}(n)+K\cdot n))$.
As outlined in Sec. 4.4, this complexity can be reduced to $\mathbf{T}(n)=\Theta(K^{\prime}(n^{2}+\Theta_{\text{mp}}(n)+K^{\prime}n))$ for the large label space, where $K^{\prime}\ll K$, denoting a significant reduction in computational cost. In practice, the entire process can be batch-processed, where the similarity is calculated among the nodes within each batch, thus making the time complexity of each batch $B$ to $\mathbf{T}(B)=\Theta(K^{\prime}(B^{2}+\Theta_{\text{mp}}(B)+K^{\prime}B))$. Computational operations of $K^{\prime}$ labels can be performed in parallel by constructing a matrix tensor for $K^{\prime}$ label spaces.
@@ A.3. Pseudocode of CorGCN
CorGCN includes (1) correlation-aware graph learning with feature and topology decompositions, and (2) correlation-aware graph convolution with intra-label message passing and inter-label correlation propagation. The general process is described in Algorithm 1.